• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name efuchey

    Log entry time 14:19:31 on September 30, 2010

    Entry number 333217

    keyword=Compton cavity status

    I started back the laser and the cavity. The cavity still works very well. The attached plot shows cavity power along time, showing that I made the cavity lock for a ~2 minutes period 6 times within 20 minutes. This is typically the "locking rate" which is used during data taking.






    Figure 1